OpenStudy (anonymous):

A mass is oscillating on a spring with a period of 1.20 s. At t = 0 the mass has zero speed and is at x = 4.85 cm. What is the magnitude of the acceleration at t = 3.30 s?

OpenStudy (anonymous):

\[F=ma=-kx\] So \[a=-\frac{k}{m}x\] Find x at t = 3.30s: \[x=A\cos(2\pi ft)\] with: f = 1/T use k/m in the above equation for 'a' from: \[f = \frac{1}{2\pi}\sqrt{\frac{k}{m}}\]

OpenStudy (anonymous):

ok,let me give it a try,thanks

OpenStudy (anonymous):

what will be my k?

OpenStudy (anonymous):

You don't know k or m, but you can solve for k/m from the last equation.

OpenStudy (anonymous):

square root of k/m is the same tin as k^2/m^2 right? trying to do the algebra

OpenStudy (anonymous):

\[\frac{k}{m} = (2\pi f)^2\]

OpenStudy (anonymous):

You square both sides, which gets rid of the square root, and leaves you with everything else squared.

OpenStudy (anonymous):

is saying my answer is wrong

OpenStudy (anonymous):

I don't have a calculator with me, so I can't check your steps. Sorry.

OpenStudy (anonymous):

can u use online calculator?

OpenStudy (anonymous):

I found one. Give me a minute to compute.

OpenStudy (anonymous):

i did (2*3.14*(1/1.20))^2

OpenStudy (anonymous):

k/m = 27.4

OpenStudy (anonymous):

the system aint accepting my answer

OpenStudy (anonymous):

x = 4.63 cm

OpenStudy (anonymous):

a = - 126.9 cm/s^2

OpenStudy (anonymous):

still wrong

OpenStudy (anonymous):

is your answer in m/s^2 rather than cm ??

OpenStudy (anonymous):

cm

OpenStudy (anonymous):

I don't see an error here. Maybe someone else can find the problem.